ISS On-Orbit Status 12 November 2001
All ISS systems continue to function nominally, except as noted previously or below. EVA-4 was completed flawlessly, 48 minutes ahead of schedule.
The crew spent most of the day with final preparations for the spacewalk this afternoon, the 29th in support of the ISS and the third from the Pirs Docking Compartment (DC-1).
FE-1 Tyurin set up SSRMS cable connections to allow video coverage of the EVA with the arm's elbow and tip LEE (latching end effector) TV cameras. Later, he calibrated the SSRMS hand controller and configured software settings.
CDR Culbertson and PLT Dezhurov configured the DC-1 and the SM Transfer Compartment (PkhO), including their communications systems, for the excursion into space. They also tested Orlan suit systems and the Orlan interface units (BSS) in DC-1 and PkhO.
On-board systems in the ISS were configured to pre-EVA status. This included deactivation of fire detection mode in the PkhO, setting of threshold limits for pressure sensors and the pressure alarm sensor (DSD), preparing the FGB and DS-1 caution and warning panels (PSS) for work during EVA, deactivation of Vozdukh, placing the Elektron O2 generator in 16 amps mode, disassembly of the air duct between SM Work Compartment (PO) and the PkhO section, disassembly of the DC-1 air duct, and others. After conclusion of the spacewalk, all ISS systems will be restored to their initial configurations tomorrow.
After a nominal regime of food and water consumption yesterday, CDR Culbertson and PLT Dezhurov today had special pre-EVA meals. Based on Russian spacewalk experience, Moscow had recommended for the CDR either a helping of grilled pork chop/pork in sour-sweet sauce, or a helping of canned fish plus some "tvorog" (cottage cheese) with nuts, for the PLT a helping of canned meat or fish and a helping of tvorog with nuts. Both were to drink apricot or peach juice before donning their Orlan spacesuits, and some hot tea or coffee, preferably sweet, after the EVA and Orlan doffing.
Both the CDR and the PLT completed standard Russian pre-Orlan-suited medical tests MO-8 (body mass measurement) and MO-9 (biochemical urine analysis). These tests will be repeated late tonight after conclusion of the spacewalk.
Culbertson and Dezhurov began suiting up at about 2:00 pm EST. After a number of further integrity checks, hatches were closed and the depressurization process in the DC-1 started between 2:35 pm and 4:05 pm. A first leak check of the hatch seal between DC-1 and PkhO failed and had to be repeated. This caused the original timeline to be delayed by about 30 minutes.
The two spacewalkers (EV1/Dezhurov red striped suit, EV2/Culbertson blue stripe) went on internal (backpack) power at 4:35 pm EST. Hatch opening was at 4:41 pm, and the CDR emerged first, on his first EVA. For Dezhurov, it was the 8th spacewalk (five were conducted from Mir). After negotiating the translation path to the SM PkhO, made a bit more difficult from this particular egress hatch by the newly installed "ladder" between the DC-1 and the SM, EV1 and EV2 mated intermodular (DC-1/SM) connectors of three low-frequency and four high-frequency cables for the KURS-P rendezvous radar system. Work at this time was already ahead by about 45 minutes.
Two additional tasks followed: survey and photography of the SM's #2 solar array wing (for which the ground commanded a slight wing rotation), and a satisfactory functional testing of the newly installed Strela cargo crane, crank-operated by EV1 from the Operator's Post mounted on the DC-1 through extension, retraction and motion in pitch and yaw direction. All tasks were successfully completed at 9:35 pm. Ingress in the airlock was 51 minutes ahead of timeline, and the hatch was closed at 9:45 pm EST, concluding the EVA after a duration of 5 hours 4 minutes. Total EVA time for Expedition 3 (Increment 3) has now reached 15h 54m, and today's spacewalk was the fifth EVA from ISS airlocks, totaling 20h 15 m). It was also the 29th EVA in support of ISS assembly, which now has run up 183h 18m of duration.
Today's optional CEO (crew earth observations) opportunities were the Eastern United States (cities under the pass: Detroit, Cleveland, Pittsburgh, and Washington/Baltimore. Multiple cities could be viewed left of track from Washington to New York and oblique photos of this area may reveal heavy aerosol loadings, but weather may prevent observation), California Central (pass over San Francisco Bay area southward into the Central Valley, but may be No-Go due to weather), High Central Andean Glaciers (synoptic shots left of track of the region with numerous volcanic peaks with permanent ice pack and snow, with small but important glacial features; again possible No-Go due to weather conditions), and the Tuamotu Archipelago (coral reef distribution: mapping strip requested with views as near nadir as possible).
Crew wakeup time on Tuesday will be at a leisurely 12:00 noon EST, and sleep begins at 4:30 pm.